OpenStudy (anonymous):

If f(x) = (x + 1)2, then f(t + 2) =

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

but their result is t2 + 6t + 9

OpenStudy (anonymous):

ok I can understand now

OpenStudy (anonymous):

thanks

OpenStudy (gw2011):

f(x)=(x+1)^2 Then: f(t+2)=(t+2+1)^2 f(t+2)=(t+3)^2 f(t+2)=(t+3)(t+3) f(t+2)=t^2+6t+9